Alan (27) and Kara (26) are married and filing a joint return. Both Alan and Kara were full-time students all year at State College, located at 11100 College Boulevard, Your City, Your State. During the year, Alan paid $5,000 in tuition and fees, substantiated with Form 1098-T. Kara paid $3,900 in tuition and fees, also substantiated with Form 1098-I. Both Alan and Kara have part-time jobs outside of school. They did not receive any scholarships during the year. Their adjusted gross income is $45,000. Their total tax on line 11 of their federal Form 1040 is $2,142. They had no additional taxes. Both Alan and Kara meet all the requirements to qualify for the American Opportunity Tax Credit.

Complete Form 8863, Education Credits (American Opportunity and Lifetime Learning Credits) for Alan and Kara.
Answer #1

As per the provisions of the law American opportunity credit allows tax reduction upto $2500 subject to condition that adjusted gross income does not exceed $90000 individually or $180000 jointly. The deduction availability is 100% of first $2000 and 25% of exceeding $2000 education expenses.

In the given case AOC available is :-

Alan : $2000 + (5000-2000×25%)= 2750 limited to $2500.

Kara : $2000 +(3900-2000×25%)= 2475.

The lifetime learning credit reduces the tax liability up to $2000 for the eligibile education expenses subject to condition that the adjusted gross income shall not exceed $68000 individually and $136000 jointly.credit is allowed 20% of education expense.

In the gives case lifetime learning credits is available is : -

Alan :- 5000×20% = 1000.

Kara :- 3900×20% = 780.